complete the ordered pairs so that each is a solution for the given equation:
3x+4y=12 (0 ), (,3/4), (, 0), (8/3,)

Answer by elima(1433)

3x+4y=12 (0 ), (,3/4), (, 0), (8/3,)
3(0)+4y=12
4y=12
y=3
======
3x+4y=12
3x+4(3/4)=12
3x+12/4=12
3x+3=10
3x=12
x=9
=============
3x+4y=12
3x+4(0)=12
3x=12
x=4
=============
3x+4y=12
3(8/3)+4y=12
24/3+4y=12
8+4y=12
4y=4
y=1
===========
:)